OpenStudy (anonymous):

Pete is standing 2 feet away from a mirror on the ground. Pete is 4 feet tall and can just see the top of the tree in the mirror from where he is standing. The mirror is 12 feet away from the base of the tree. How tall is the tree? 6 feet 24 feet 14 feet 22 feet

hartnn (hartnn):

draw a figure...

hartnn (hartnn):

just use proportionality

hartnn (hartnn):

hmm, not 20, 2/12 = 4/x

hartnn (hartnn):

i used the fact that size of reflection S is directly proportional to distance D S1/D1 = S2/D2

hartnn (hartnn):

i'll keep them, if someone google searches this question, he/she will get an explanation